OpenStudy (anonymous):

Help me pleaze : Resolve the equation in IR: cos^5(x)-sin^5(x)=1

OpenStudy (anonymous):

Notice that the \[ 2 k \pi \text { for } k=0, \pm1, \pm2, \pm 3, \cdots\\ and \\ -\frac \pi 2 + 2 k \pi \text { for } k=0, \pm1, \pm2, \pm 3, \cdots\\ \] are solutions.

OpenStudy (anonymous):

How do you find it ? pleaz

OpenStudy (anonymous):

For what x does sin(x)=0 and cos(x)=1? For what x does cos(x)=0 and sin(x)=-1?

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

How did you find that sinx = 0 or cosx = 0

OpenStudy (anonymous):

If sin(x)=0 and cos(x)=1 then cos^5(x) - sin^5(x)= 1-0=1 If sin(x)=-1 and cos(x)=0 then cos^5(x) - sin^5(x)= 0-(1)^5=0+1=1

OpenStudy (anonymous):

These are may be some of the solutions, but if you use a graphing method, then you realize that these are the only one

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

I have another problem can u help me

OpenStudy (anonymous):

\[(a,b,c)\in IR ^{3}: abc=1 prove that: (a-1+\frac{ 1 }{ b })*(b-1+\frac{ 1 }{ c })(c-1+\frac{ 1 }{ a })\le 1\]

OpenStudy (anonymous):

Replace c by 1/(ab) in the expression to get \[ \left(a+\frac{1}{b}-1\right) \left(\frac{1}{a}+c-1\right) \left(b+\frac{1}{c}-1\right)= \\ \left(a+\frac{1}{b}-1\right) \left(\frac{1}{a b}+\frac{1}{a}-1\right) (a b+b-1) \] Call \[ f(a,b)=\left(a+\frac{1}{b}-1\right) \left(\frac{1}{a b}+\frac{1}{a}-1\right) (a b+b-1)\\ \text{ Notice that } \\ f(1,1)=1 \]

OpenStudy (anonymous):

The way a I did it, I found that the gradient of f is zero at (1,1). The Hessian determinant of f at (1,1) is 3>0 and the second derivative with respect to a at (1,1) is -2. Hence f has a maximum at (1,1) which is 1. We are done. The computations should be done by a machine and that is how I did them and I refuse to do them by hand.